Operations Report April 25th

Crew 193 Operations Report 25Apr2018

SOL: 3

Name of person filing report: David Attig

Non-nominal systems: Com radio #6, Sim suits #10, 2, 5, 4, 6.

Notes on non-nominal systems: Radio #06 (newer style) only lasted 1 hour on EVA with a freshly charged battery. Radio will not be used for future EVAs. Sim suit 10 still has the badly scratched visor, suits 2, 5, 6 need maintenance on straps attaching neck ring to pack, and suit 4 has a broken charge connector.

Generator (time run): 2000-0715

Solar— SOC 95 % (Before generator is run at night)

Diesel – 75 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 3500 gallons.

Water (storage) – 550 gallons

Water (static) – 350 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134078.1

Toilet tank emptied: no
Date last emptied: April 24

ATV’s (Honda, 300, 350.1, 350.2, 350.3):
Honda Used?: Yes
300 Used? No
350.1 Used? Yes
350.2 Used? No
350.3 Used? No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 1/4 Gallons

# Hours the ATVs were Used today: 1.2

Notes on ATVs: None

Deimos rover used: No
Hours: 124.0
Beginning charge: 100%
Ending charge: 100 %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 33.7
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 23
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: We would like the sim suit supply box dropped off in our airlock, so we can work on the two-piece suits.

Summary of internet: Appears to be working.

Summary of suits and radios: Radio 06 appears to transmit a beep continuously, draining the battery. Suits have some issues; see email titled “Details about EVA Suits”.

Summary of Hab operations: supported two EVAs and investigated suit status. Sol Summary report for more details.

Summary of GreenHab operations: plants are doing well, see GreenHab report for more details.

Summary of ScienceDome operations: Progress was made on dishwasher experiment, see Science Report for more details.

Summary of RAM operations: Placed tools in RAM for future use, took boards apart and painted some to use for an MDRS sign.

Summary of health and safety issues: Need to work on better balance between working in RAM and not breaking sim; see email “A few notes – about EVA duration” for details.

Questions, concerns and requests to Mission Support: A few requests, listed in Sol Summary.

EVA Report April 25th

Crew 193 EVA Report 25Apr2018
Sol 2 Crew 193 EVA 6 Summary
Earth-Date 25APR2018

EVA Gold Crew Members – Gold001 (XO Campbell) / Gold004 (HSO Pandya) (Informally: “Team Canada”)

EVA Objectives – Summary

1. Exploration of The Moon – Could not adequately confirm en route; map now suggests we overshot and ended up in Lith Canyon. Suggest re-visiting actual Moon site at a later date.

Distance covered by foot – 11,500
Steps – 8.5 km

2. Scouting and Mapping of Emergency Shelter Sites – Not Mapped, as uncertain regarding location / Raw data here

3. EVA Comfort, Well-Being, Safety and Optimization Study / Raw data here

4. Visor Fogging Study – Data collected / Raw data here

Gold001’s comms failed almost immediately flashing ‘low battery’ despite having been charged x 24 hours. This was mitigated intra-EVA by maintaining line-of-sight contact at all times, and turning on this unit only when LOS was not possible.
Were unable to tell where the turn-off for White-Yellow-Beige Moon region was; suspect we overshot and ended up in Lith Canyon. Will review images and maps to confirm.

ATVs Used

Manual gear-shift (red)
Blue ATV #21
These ATVs require fuel (requested); will refuel before next use.

Safety Measures & Preparation

Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided Adequate hydration – No
Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

Sol 2 Crew 193 EVA 7 Summary
Earth-Date 25APR2018

EVA Gold Crew Members – Gold000 (CO Patil-Sabale) / Gold002 (CE Attig) / Gold005 (CGA Shear)

EVA Objectives – Summary

Paint the MDRS Signpost in the RAM
Climb the hills behind the hab to head to the Hab Ridge, explore Hab Ridge to locate emergency shelter/s

Painting in the afternoon hours in the extremely hot RAM proved unhealthy. After 45-50 minutes in the RAM, the crew started trekking up a hill to get to the Hab Ridge. Halfway up the hill our CO felt dehydrated. The EVA was cut short and the crew returned to the RAM. Need fans in the RAM
Discontinue using prototype suit for EVA’s during this mission, it’s too small and doesn’t allow enough airflow which is a must especially during a hot day.

ATVs Used
None. This was a planned project and on-foot EVA

Safety Measures & Preparation

Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided Adequate hydration – Yes
Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

Sol Summary Report April 25th

Crew 193 Sol Summary Report – 25 APR 2018

Sol 3

Summary Title: Life Will Find a Way

Author: Doug Campbell, Executive Officer and Science Officer

Mission Status: Sub-optimal due to minor crew health issues but expected to be back to nominal tomorrow

Sol Activity Summary: The day started off early with quick individual breakfasts in order to prepare for EVA 6 which left at 0730. EVA 6 was a successful failure in that valuable exploration of Lith Canyon was completed; however, the intended destination was the White Moon area! While EVA 6 was occurring, the plants were watered in the Green Hab and some vegetables were harvested. Our crew astronomer made a quick trip to the observatory and did some troubleshooting with the GPS units. Our crew engineer did work around the hab and communicated with mission control.

The crew lunched together (thanks to our commander for cooking!) and then spent time on personal work and research projects before EVA 7. EVA seven was cut short at 1600 by the Health and Safety Officer over concerns of crew dehydration. Late afternoon was spent on reports. The plan for tonight is to watch a movie as a team, perhaps…. The Martian!!

Look Ahead Plan: EVA 8 will attempt to complete the mission plan that was submitted for EVA 6 but not completed. EVA 9 will be exploring further into Candor Chasma as part of the GPS mapping study. We will continue to work on learning ASL for 5 minutes each day as well as sharing stories about our lives so that the crew gets to know each other better.

Anomalies in work:

1. Radio #6 has a poor battery which loses charge quickly. It caused Gold 001 to be without communication during EVA 6. This radio has been put out of service by crew 193
2. EVA 7 was cut short due to health concerns over crew dehydration

Weather: Clear skies in AM, partly cloudy and gusty in PM.

Crew Physical Status: Some blisters, one incidence of crew mild dehydration. All members are resting and back to nominal by days end.

EVA 6 – Successfully explored the edges and the base of Lith Canyon. As per the title of this report, the EVA team found group of fully blossomed wild flowers on the edge of a large, sandy rock formation with no other flora or fauna in sight; proving that even on mars, life will find a way.
EVA 7 – Began the EVA in the RAM to paint the new sign for the MDRS campus. The crew then proceeded to climb the hill up to hab ridge. However, during the climb, one of the members began to not feel well and the remainder of the EVA was scrubbed.

Reports filed:
Science Report
Astronomy Report
Journalist’s Report
Photo & Photo of the Day Report

Reports to be filed:
Operations Report
EVA Report

Support Requested:

Additional fuel for ATVs.
Orange or green paint for the MDRS sign (or both)
Hot pads so we can put hot cookware on the table without wrecking the tabletop.

Science Report April 25th

Hello Mission Support,

Here is the Science Report for 25 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by CGA Eric Shear

1. Emergency EVAC EVA: EVA 6 did not find any suitable locations for evacuation in Lith Canyon. The distance is very far from the hab, difficult to reach and quite exposed to the environment. Although it is beautiful! EVA 7 was to walk south alongside the Hab Ridge Road on foot, but was cut short due to warm conditions. Further investigation of possible locations will be done on future EVAs.

2. Sunspot and Solar Flare Monitoring: The sunspot from the previous observation has split into three smaller ones, and one prominence was seen. For photos, please refer to the astronomy report.

3. Shortwave Texting on EVA’s: After troubleshooting, one Beartooth device continues to not work today. We are awaiting tech support from the manufacturer. In the meantime, we continue to use the note app on our phones to communicate with our deaf crew member.

4. GPS Route Measurements: A GPS tracking route was successfully taken on EVA 7 with the Garmin GPSmap 60Cx. On this EVA, we walked 641 meters (roundtrip).

5. Waterless dish cleaning: Baseline testing was completed this afternoon. 2 swabs were taken from dirty dishes in the sink and two swabs were taken from dishes in the cupboard. The swabs were rubbed on a water/agor mix in petri dishes and will be watched over the next few days for bacteria growth.

Astronomy Report April 25th

Astronomy Report
Name: Eric Shear
Crew: 193
Date: April 25, 2018

Sky Conditions: Clear

Wind Conditions: None

Observation Start Time: 08.30

Observation End Time: 09.15


Our CGA continued his task of monitoring the sun for sunspot activity. The sunspot observed yesterday has now split into three smaller ones. One larger prominence was visible with the eye, but not with the Skyris camera. No flares or filaments were seen.

Photos were taken with the Skyris camera and the best image processed in Autostakkert, Registax, Photoshop, and PowerPoint to colorize and bring out contrast. The results are attached. The image labeled “prominence” is where the prominence was seen, but the camera did not pick up on it.

Objects Viewed: Sun

Problems Encountered: Skyris camera did not show the prominence that was visible with the eye, no matter how the image was processed.

Journalist Report April 25th

Journalist Report – Earth-Date 25APR2018 / Sol 3
Author: Shawna Pandya

Never a dull moment on Mars! Over the past 24 hours, this intrepid and resilient crew has bonded over learning the basics of American Sign Language (fun fact: you can actually count to 999 on one hand) and sharing stories of diving with sharks, meetings astronauts, driving in Europe, chasing typhoons in Guam and scaling peaks in the Rockies (told you we weren’t dull!).

Our day today started with a rude wake-up at 0300 local time as the Commander and HSO went to investigate a persistent banging in the rear airlock (and the HSO learned a valuable life-lesson about
double-checking the security of the bolt on the outer airlock door in future instances).

The day continued with a set of adventures, as EVA Team 1 (Team Canada!) overshot their intended scouting site, landing in Lith Canyon instead of the Moon – luckily some good scientific data was still accrued, and the Team will attempt to shoot for the Moon again tomorrow. EVA Team 2 furthered their progress with the MDRS sign project in the RAM during their mid-day EVA, but responsibly chose to cut their EVA short some time later, realizing the real risk of heat stroke from working in the hot RAM at mid-Sol.

Science, GreenHab and Hab operations continued on as usual, and Gold Crew will cap the Sol off with an evening of R&R that includes a movie, some more team-bonding through shared adventures and ASL lessons, and a delicious Martian spread of spaghetti alfredo, spaghetti with Martian tomato sauce, and some good ol’ fashioned Hab-made cornbread.

Gold Crew 193 looks forward to greeting Sol 4 with enthusiasm, and facing any challenges the Martian winds bring the same way we always have – as a team, and with enthusiasm!

Copyright © The Mars Society. All rights reserved. | Main Site